OpenStudy (anonymous):

please solve.. f(x)= 3x+4 f(3), f(-2), f(-5).. :)..

OpenStudy (mayankdevnani):

\[\huge f(x)=3x+4\] so, \[\huge f(3)=3(3)+4\] \[\huge f(-2)=3(-2)+4\] \[\huge f(-5)=3(-5)+4\]

OpenStudy (mayankdevnani):

can you solve it??? @mjay

OpenStudy (anonymous):

yes..

OpenStudy (mayankdevnani):

good....so tell me answers

OpenStudy (anonymous):

wait i will solve it..

OpenStudy (mayankdevnani):

ok...

OpenStudy (anonymous):

f(x)=x+7 f(3)=3+7 f(3)= 10 answer.. f(-2)= -2+7 f(-2)= 5 answer.. f(-5)= -5+7 f(-5)= 2 answer..

OpenStudy (mayankdevnani):

you are wrong!! @mjay

OpenStudy (anonymous):

oh... what's wrong??

OpenStudy (mayankdevnani):